OpenStudy (anonymous):

Write the standard equation for the circle center (–6, 7), r = 9. (1 point) (x – 6)2 + (y + 7)2 = 9 (x – 7)2 + (y + 6)2 = 81 (x + 6)2 + (y – 7)2 = 81 (x + 6)2 + (y – 7)2 = 9

OpenStudy (anonymous):

. Write the standard equation for the circle center (–6, –8), that passes through (0, 0). (1 point) (x – 6)2 + (y – 8)2 = 10 (x – 6)2 + (y – 8)2 = 196 (x + 6)2 + (y + 8)2 = 14 (x + 6)2 + (y + 8)2 = 100

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2 where h and k are the coordinates (h,k) which would give you (x+6)^2+(y-7)^2=81 for the first one..

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

no